OpenStudy (anonymous):

What value of x makes the ratios equivalent? eight over x equals ten over fifteen

OpenStudy (anonymous):

\[8/x=10/15\] \[8*15/10=x\] \[x=12\]

OpenStudy (anonymous):

Thank you so much now i understand this

OpenStudy (anonymous):

Every fraction = to another fration can be done the same way for the value of x.